OpenStudy (anonymous):

Percent change from 24 to 72

jimthompson5910 (jim_thompson5910):

First subtract: 72 - 24 = ???

OpenStudy (anonymous):

48

jimthompson5910 (jim_thompson5910):

Now divide this by the first value 24 48/24 = 2 Then multiply by 100 to convert to a percentage: 2*100 = 200%

jimthompson5910 (jim_thompson5910):

so the percent change from 24 to 72 is 200%

OpenStudy (nathan917):

Percent change= change in value/ original value X 100 So... percent change= 72-24/24 X 100 Percent change= 200% It's a positive number so it must be an increase Choice D. 200% increase

jimthompson5910 (jim_thompson5910):

how do you know it's choice D???

jimthompson5910 (jim_thompson5910):

no answer choices were given

OpenStudy (nathan917):

done this test...